OpenStudy (anonymous):

For $7.52, you purchased 8 pens and highlighters from a local bookstore. Each highlighter costs $1.09 and each pen costs $.69. How many pens did you buy. I've gotten down 1.09h + .69p = 7.52 but that's as far as I've gotten. If anyone could help me work this

OpenStudy (anonymous):

i can. it says you got 8 pens. 1.09h + .69(8) = 7.52 1.09h + 5.52 = 7.52 do you know what to do from here?

OpenStudy (anonymous):

I thought that the 8 was for the combines amount of pens and highlighters?

OpenStudy (anonymous):

This is for a study guide and the other questions in the lesson were about systems of equations, so because of the format of those, I'm completely lost when it comes to this problem.

OpenStudy (anonymous):

no. it say 8 pens and highlighters. i guess it seems confusing.. but i don't think they would do that. to me it seems 8 pens, and some highlighters. the other way it wouldn't make sense. do you see? can you show me another one in the guide, though, so i can see?

OpenStudy (javeneno):

I think that it a total of 8 pens and highlighters

OpenStudy (javeneno):

because if it was 8 pens and some highlighters it would say that

OpenStudy (anonymous):

Do you mean just the other problems in general or word problems, because that was the only word problem.

OpenStudy (javeneno):

the word problem

OpenStudy (anonymous):

These are the other problems, and I got kind of confused from the earlier examples in the book.

OpenStudy (javeneno):

im doing number 9

OpenStudy (anonymous):

yes, that's the one

OpenStudy (anonymous):

then i'm not sure... sorry ^^

OpenStudy (javeneno):

I think i got it. Silverr was right it is 8 pen and some highlighters

OpenStudy (javeneno):

You have to multiply 8 x $1.09 (price)

OpenStudy (anonymous):

So the first equation is 1.09h + .69(8) = 7.52?

OpenStudy (javeneno):

yes

OpenStudy (anonymous):

okay, than :) 1.09h + .69(8) = 7.52 1.09h + 5.52 = 7.52

OpenStudy (anonymous):

ok whats next

OpenStudy (anonymous):

then it would be 1.09h = 2?

OpenStudy (anonymous):

1.09h + 5.52 = 7.52 7.52 - 5.52 = ?

OpenStudy (javeneno):

so subtract 5.52 from 7.52 which is 2.00. so he has $2.00 left how many highlighters can he get?

OpenStudy (anonymous):

Just 1?

OpenStudy (javeneno):

yes

OpenStudy (javeneno):

so the answer should be 1

OpenStudy (anonymous):

I'm still confused on the end of the question though, because it asks how many pens you bought

OpenStudy (javeneno):

Im verry sorry if i confused you but it is just 8 pens and highlighters. (:-/

OpenStudy (javeneno):

i actually dont know how to solve it but your supposed to find the number of pens you can still afford with the highlighters

OpenStudy (anonymous):

Oh, that's alright. Thank you so much for your help!

OpenStudy (javeneno):

@phi

OpenStudy (javeneno):

maybe phi can help ,

OpenStudy (anonymous):

Ok, thanks again

OpenStudy (phi):

It is poorly written, but they probably mean h+p= 8 use that along with 1.09h + .69p = 7.52 to solve for h and p

OpenStudy (anonymous):

I hope it's not too much to ask, but could you help walk me through that?

OpenStudy (anonymous):

I feel like I have to move one of the variables to the other side with 7.52 to solve it?

OpenStudy (phi):

you want to solve for p, so I would multiply the first equation h+p= 8 by -1.09 (that means both sides , all terms) then add the two equations. (this is using "elimination") you said I feel like I have to move one of the variables to the other side with 7.52 to solve it? yes, you can take that approach (called "substitution") But I would get used to elimination... it works better for hairier problems.

OpenStudy (anonymous):

yeah, I tend to struggle a lot with the elimination method. So it would be 1.09h + 1.09p = 8.72 for the second equation right?

OpenStudy (phi):

you should get -1.09h + -1.09p= -8.72 1.09h + . 69p = 7.52 --------------------- now add them

OpenStudy (anonymous):

where did the negative terms come from on the top equation?

OpenStudy (phi):

I multiplied by -1.09 so that when we add the "h" terms cancels

OpenStudy (phi):

start with h+p=8 multiply by -1.09

OpenStudy (anonymous):

So every time you do that you multiply with a negative term?

OpenStudy (anonymous):

-.4p = -1.2

OpenStudy (anonymous):

so p would equal 3?

OpenStudy (phi):

I looked at the second equation 1.09h + . 69p = 7.52 and asked, "what do I add to 1.09h to make it go away? " answer: -1.09h then I looked at h+p=8 and asked, what do I multiply h by to get -1.09 h ? answer: -1.09 so multiply the h+p=8 equation by -1.09 all terms, both sides to get -1.09 h + -1.09p = -1.09*8

OpenStudy (anonymous):

oh, ok that makes sense now

OpenStudy (phi):

yes looks good. p= -1.2/-0.4 =3

OpenStudy (anonymous):

does it matter which equation I substitute p into?

OpenStudy (anonymous):

when I did it with h + p = 8, I got h =5

OpenStudy (phi):

to find h ? no. But one looks easier to use than the other i.e. h=5 but both equations will work.

OpenStudy (phi):

though it's not obvious, the two equations can be thought of as equations of lines, and the solution is the point on both lines. So both equations will give the same answer for h, if you use p=3

OpenStudy (anonymous):

I think I get it now. Thank you so much!

OpenStudy (phi):

example 1.09h + . 69p = 7.52 1.09h + . 69*3 = 7.52 1.09h + 2.07 = 7.52 1.09h = 7.52-2.07 1.09h= 5.45 h= 5.45/1.07 = 5

OpenStudy (phi):

obviously the same answer, but a lot more work than 3+h= 8 h= 5